OpenStudy (anonymous):

What is the solution to the equation?    5(8 + 7) = 5 • x + 5 • 7

OpenStudy (anonymous):

First you would distribute the 5 in "5(8 +7)"

OpenStudy (anonymous):

Giving you this \[40 + 35 = 5 \times x + 5 \times 7\] Then you would gather your like terms on both sides.